# Balanced Resonate-and-Fire Neurons

### v1

@@ 7 Discussion
The results from our BRF-RSNN implementation lead to considerable insight into the workings of the RF neurons. Importantly, the learning curve and parameter analyses of the datasets show that the BRF networks learned meaningful parameters by favorably tuning the resonating behavior, modeling the complex dynamics not yet observed in large-scale simulations. We observe favorable oscillatory behavior of the membrane potential and increasing frequency of periodic spikes with higher angular frequency in the raster plot for all datasets. For SHD, output spikes are present towards the end of the sequence, even with the zero padding of the input signal. As most signals are still periodic, it suggests controlled oscillation and spiking that maintained relevant frequencies until the end of the sequence.
Note that we conducted the experiments without biases for all RF models to see how many spikes are needed to keep the network dynamics alive and to solve the tasks. As a result, a few of the BRF neurons in the network learned to fire consistently throughout the sequence as seen in Figure 6 and thus “imitated” a bias injecting constant current into the network. In contrast, ALIF networks need explicit biases to perform well. The BRF networks are thus more flexible in modeling behavior that cannot be learned by the integrator neurons alone.
@@ 8 Conclusion
We introduced BRF- and BHRF-RSNNs as a principled resonating spiking neuron model that solves the training difficulties encountered for networks comprised of standard RF neurons while demonstrating the effectiveness of internal resonating state as a form of long-term memory. Our BRF- and BHRF-RSNNs with smaller network architectures and sparser spiking outperformed and yielded results comparable to the deep baseline models. The learning curves showed that the BRF models converged faster than the ALIF models. Furthermore, the stable convergence hints at good reproducibility of the BRF model given the hyperparameters. The BRF parameter analyses displayed indications of the network to learn frequencies underlying the input signals and showed resilience against sparser recurrent connections compared to the ALIF network.
One drawback of our BRF- and BHRF-RSNNs was the difficulty in finding an optimal initial parameterization of the angular frequencies and the dampening factor offset, since both determine the models’ performance, The analyses of the parameters introduced the possibility of applying the Fourier transform on the datasets a-priori to approximate the range of angular frequencies the BRF neurons requires to optimally learn the data.
Like ALIF networks, the BRF-RSNNs required extensive time for training due to the custom implementation in PyTorch; one training period of the S-MNIST and PS-MNIST, for example, took 20 to 30 hours with a NVIDIA GeForce RTX 2080 Ti, similar to the ALIF single-layered model. The training period was shorter compared to the deep-layered ALIF networks, but still remains a challenge. Nonetheless as shown in the results, faster convergence enabled earlier stopping of training with acceptable performance if a short training period was required, for example, in practical applications.
Possible future research may be to scale up the simulations to much larger and more complex problems, such as the CIFAR (Krizhevsky & Hinton, 2009) or the Google Speech Command dataset (Warden, 2018). It is of particular interest to apply the BRF model to raw audio processing task, because of its ability to extract periodic pattern within the time domain.Moreover, applying trainable refractory period or completely adaptive threshold may be fascinating.
This study was an initial exploration to investigate the working of resonating neurons in the recurrent spiking neuron framework. Our networks showed comparable results and exceeded the state-of-the-art for RSNNs. We see many possible approaches to extending our work, as the implementation was kept to its simplest form. The presented work thus provided a foundation for further exploration of the BRF-RSNN variants.
@@ A.1 Traditional soft and hard reset
The traditional reset mechanism considered was formulated as the following:
$u(t)=(1-\zeta\,z(t))\,u^{\prime}(t)$ (23)
for which the $u^{\prime}(t)$ represent the membrane potential before reset and $\zeta\in(0,1]$ the amount of reset, where $\zeta=1$ led to a hard reset, otherwise soft.
@@ A.7 ECG-QT database preprocessing
A.8 Hyperparameters
Table 3: Sequential-MNIST hyperparameters applied for the best-performing BRF, HBRF, and ALIF model with label-last loss and truncation step of 50 for TBPTT (used to compare pruning). $\mathcal{U}(a,b)$: parameters sampled from a uniform distribution with range of $a$ to $b$. $\mathcal{N}(\mu,\sigma)$: parameters sampled from a normal distribution with mean $\mu$ and std $\sigma$.
Table 4: Permuted S-MNIST hyperparameters applied for the best-performing BRF, HBRF, and ALIF model with average sequence loss and truncation step of 50 for TBPTT.
Model BRF-RSNN BHRF-RSNN ALIF-RSNN
BPTT
TBPTT
BPTT
BPTT
TBPTT
@@ A.8 Hyperparameters
$\omega:\mathcal{U}(15,50)$
$b^{\prime}:\mathcal{U}(0.1,1)$
@@
$\omega:\mathcal{U}(15,35)$
$b^{\prime}:\mathcal{U}(0.1,1)$
@@
Model BRF-RSNN BHRF-RSNN ALIF-RSNN
BPTT
TBPTT
BPTT
BPTT
TBPTT
@@
$\omega:\mathcal{U}(15,85)$
$b^{\prime}:\mathcal{U}(0.1,1)$
@@
$\omega:\mathcal{U}(15,50)$
$b^{\prime}:\mathcal{U}(1,6)$
@@
Table 4: Permuted S-MNIST hyperparameters applied for the best-performing BRF, HBRF, and ALIF model with average sequence loss and truncation step of 50 for TBPTT.
Table 5: ECG QT database hyperparameters applied for the best-performing BRF, HBRF, and ALIF model with truncation step of 50 for TBPTT.
Table 6: Spiking Heidelberg dataset (SHD) hyperparameters applied for the best-performing BRF, HBRF, and ALIF model with average sequence loss and truncation step of 50 for TBPTT.
Model BRF-RSNN BHRF-RSNN ALIF-RSNN
BPTT
TBPTT
BPTT
BPTT
TBPTT
@@
$\omega:\mathcal{U}(3,5)$
$b^{\prime}:\mathcal{U}(0.1,1)$
@@
$\omega:\mathcal{U}(7,11)$
$b^{\prime}:\mathcal{U}(0.1,1)$
@@
Model BRF-RSNN BHRF-RSNN ALIF-RSNN
BPTT
TBPTT
BPTT
BPTT
TBPTT
@@
$\omega:\mathcal{U}(5,10)$
$b^{\prime}:\mathcal{U}(2,3)$
@@
$\omega:\mathcal{U}(3,10)$
$b^{\prime}:\mathcal{U}(0.1,1)$
@@
Table 6: Spiking Heidelberg dataset (SHD) hyperparameters applied for the best-performing BRF, HBRF, and ALIF model with average sequence loss and truncation step of 50 for TBPTT.

### v2

@@ 7 Discussion
The results from our BRF-RSNN implementation lead to considerable insight into the workings of the RF neurons. Importantly, the learning curve and parameter analyses of the datasets show that the BRF networks learn meaningful parameters by favorably tuning the resonating behavior, modeling the complex dynamics not yet observed in large-scale simulations. We observe favorable oscillatory behavior of the membrane potential and increasing frequency of periodic spikes with higher angular frequency in the raster plot for all datasets. For SHD, output spikes are present towards the end of the sequence, even with the zero padding of the input signal. As most signals are still periodic, it suggests controlled oscillation and spiking that maintains relevant frequencies until the end of the sequence.
Note that we conducted the experiments without biases for all RF models to investigate how many spikes are needed to keep the network dynamics alive and to solve the tasks. As a result, a few of the BRF neurons in the network learned to fire consistently throughout the sequence as seen in Figure 6 and thus effectively “imitated” a bias injecting constant current into the network. In contrast, ALIF networks need explicit biases to perform well. The BRF networks are thus more flexible in modeling behavior that cannot be learned by the integrator neurons alone.
The neurons learn their own favored frequencies and, consequently, focus on individual time scales. This can be loosely seen as a homeostatic regulatory process maximizing local information gain in the realms of distributed, collective task solving within the recurrent networks.
Further exploration of the post-trained models against various noise types and numerical quantization (Stromatias et al., 2015; Park et al., 2021), shows consistently high robustness of the BRF networks compared to the ALIF networks, detailed in Section A.10, Figure 12. The BRF networks are slightly more robust than the RF networks for some noise variants, despite the constant low SOP of the BRF networks. Hardware implementations are prone to such noises (Stromatias et al., 2015), which suggests the BRF neurons to be suitable for hardware applications. We further aim to realize inference with these neurons on Loihi 2, which provides support for RF neuron variations (Davies et al., 2018; Shrestha et al., 2023).
Concerning the network dimension, we found that adding additional resonator neurons into the single hidden layer beyond a certain capacity did not significantly improve the performance further.
One drawback of our BRF- and BHRF-RSNNs is the difficulty in finding an optimal initial parameterization of the angular frequencies and the dampening factor offset, since both determine the models’ performance. The analyses of the parameters suggest the possibility of applying the Fourier transform on the datasets a-priori to approximate the range of angular frequencies the BRF neurons require to optimally learn the data.
It should be noted that the computational complexity of BRF-RSNNs is similar to single-layered ALIF models. However, due to the faster convergence the training can be stopped much earlier, which effectively results in a drastically reduced training time. Implementation-wise, another speed-up can be achieved by using forward gradient injection (Otte, 2024) for modeling the surrogate gradient functions in combination with automatic model optimization routines such as TorchScript in PyTorch (some results are presented in Section A.13).
@@ 8 Conclusion
We introduced BRF- and BHRF-RSNNs as a principled resonating spiking neuron model that solves the training difficulties encountered for networks comprised of standard RF neurons while demonstrating the effectiveness of internal resonating state as a form of long-term memory. Our BRF- and BHRF-RSNNs with smaller network architectures and sparser spiking outperformed deep baseline models. The learning curves showed that the BRF networks converged much faster than the ALIF networks. Furthermore, the stable convergence hints at good reproducibility of the BRF model’s performance. The BRF parameter analyses indicate the network to learn frequencies underlying the input signals and showed resilience against sparser recurrent connectivity compared to the ALIF network.
Possible future research may be to scale up the simulations to much larger and more complex problems, such as the CIFAR (Krizhevsky & Hinton, 2009) or the Google Speech Command dataset (Warden, 2018). It is of particular interest to apply the BRF model to raw audio processing tasks, because of its ability to extract periodic pattern within the time domain.
Another line of research could be to combine BRF neurons with online learning approaches, such as e-prop (Bellec et al., 2020), which would broaden its range of applications due to memory efficiency and faster processing of longer sequences. Additionally, trainable refractory period decays for flexible adaptive thresholding could be introduced, as well as trainable simulated variable time constants to foster multi-temporal resolution processing.
This study marks an initial exploration to investigate the working of stable resonating neurons in the context of recurrent spiking neural network. Our models show comparable results and/or exceeds the state-of-the-art for RSNNs. We see many possible approaches to extending our work, as we kept the implementation to its simplest form. The presented work thus provides a foundation for further research of possible BRF-RSNN variants.
@@ Acknowledgements
Impact Statement
The aim of this work is to advance the field of machine learning, with a specific focus on spiking neural networks (SNNs) to augment the effectiveness and energy efficiency of AI systems, making them more sustainable. It is worth noting that, like most developments in machine learning research, our work may have various societal consequences, but none of which we think needs special consideration here.
@@ A.1 Traditional soft and hard reset
The traditional soft and hard reset mechanisms were formulated respectively:
$u(t)=u^{\prime}(t)-(1+i)\,z(t)\,\vartheta$ (23)
$u(t)=[1-(1+i)\,z(t)]\,u^{\prime}(t)$ (24)
for which $u^{\prime}(t)$ represents the membrane potential before reset.
@@ A.7 ECG-QT database preprocessing
A.8 Effect of refractory period and smooth reset
Refractory period (RP) and smooth reset (SmR) both result in reduced spiking operations, with their effects varying depending on the dataset. The SOP of the BRF neuron consistently yields the lowest SOP. It suggests the flexibility of the BRF neuron to learn various types of data efficiently, compared to the RF neuron. ECG optimized with RP spikes less than with SmR, whereas SHD spikes more with SmR; however even more efficient in both cases is the BRF with the combination of RP and SmR. PS-MNIST diverges with: no reset, RP, SmR, and combination of RP and SmR. Thus, it shows significantly high SOP for these variants. It also highlights the stability brought out by the divergence boundary (DB). S-MNIST trained with RP results in unstable performance and SOP.
Figure 10: SOP result in RF networks with divergence boundary (DB), refractory period (RP), and smooth reset (SmR). Mean SOP over 5 runs plotted with standard deviation. DB, RP & SmR denote the complete BRF neuron.
A.9 Reset variation and performance
To explore the impact of the standard reset mechanisms (mentioned in Section A.1) on the resonator network, RF models with smooth, soft, and hard reset are optimized for ECG and SHD, as shown in Figure 11. Note that the traditional reset mechanisms applied to S-MNIST and PS-MNIST failed to converge. The figures show performance drop and slower convergence with the soft and hard reset, presumably due to the altered phase of the oscillation and the difficulty in continuous resonant spiking, as spikes may occur with twice the period of the original signal. Further facilitated by the result without reset, in which the convergences are on-par with the BRF neurons where the phase itself is preserved. As mentioned in the main text, the system did not diverge for the RF without reset for ECG and SHD in particular, due to the small angular frequencies leading to smaller error accumulation over time.
Figure 11: Convergence comparison of ECG and SHD over RF reset mechanisms.
A.10 Noise robustness
Noise robustness of post-trained models are further investigated with the methods applied by Stromatias et al. (2015) and Park et al. (2021). Hardware implementations of SNNs pose limitations to the performance due to restricted precision or noise in the signal. Here, four different type of restrictions and noise are explored. Quantization reduces the bit precision representation (Figure 12 top left). Noise in the input signal from the dataset are simulated with Gaussian noise with increasing standard deviation (Figure 12 top right). For spike deletion (Park et al., 2021) level, a percentage of the spikes are removed from the network at each sequence step (Figure 12 bottom left). Synaptic noise added Gaussian noise to the weights.
Figure 12: S-MNIST noise robustness studies conducted on saved best performing models. Solid line denote performance and dotted line the spiking operations (SOP) for each network.
Overall the vanilla RF network without reset and the BRF network performed better when faced with various noise than the ALIF network. Especially for the noise introduced in the input, the RF and BRF network maintained their performance, while the ALIF network failed already with a small standard deviation in noise. For the BRF network, this is achieved despite the high spiking sparsity of the BRF network compared to the RF or ALIF network. The RF network is more robust against spike deletion than the BRF or ALIF network. This may be an effect of redundant spiking for the RF networks, which highlights a trade-off of spike deletion robustness and high spiking sparsity.
A.11 Error landscape
BRF-RSNN landscape has smooth and convex-like structure, whereas landscape of RF- and especially, ALIF-RSNN, has a rough surface with a narrow valley. Such smooth landscape represent high generalization and straightforward optimization, accounting for the fast convergence.
We further found that the spectral radius of the membrane state transition mapping computed over one discrete time step results in unity or below by implementing the divergence boundary, thus effectively stabilizing the gradient and preserving its magnitude (Higuchi et al., 2024). This indicates that the divergence boundary contributes to fast and smooth learning.
Figure 13: Error landscape plots for RF, BRF, and ALIF network on the S-MNIST dataset adopted from Higuchi et al. (2024). Top: error surface plots. x and y axis correspond to $\alpha$ and $\beta$, i.e. parameter deviations, and the z axis to $f(\alpha,\beta)$, i.e. error (Li et al., 2018). Bottom: error contour plots. Note that the value range and corresponding coloring differ between the diagrams to enhance visualization.
A.12 Hyperparameters
Table 3: Hyperparameters applied for the best-performing BRF, HBRF, and ALIF model with label-last loss (S-MNIST) and average sequence loss (PS-MNIST, ECG, SHD) and truncation step of 50 for TBPTT (in pruning). Same hyperparameters applied for BRF- and RF-BPTT models, where $b=b^{\prime}$ for RF-RSNN.
S-MNIST BRF-RSNN BHRF-RSNN ALIF-RSNN
BPTT TBPTT BPTT BPTT TBPTT
@@ A.12 Hyperparameters
$\omega:\mathcal{U}(15,50)$, $b^{\prime}:\mathcal{U}(0.1,1)$,
@@
$\omega:\mathcal{U}(15,35)$, $b^{\prime}:\mathcal{U}(0.1,1)$,
@@
PS-MNIST BRF-RSNN BHRF-RSNN ALIF-RSNN
BPTT TBPTT BPTT BPTT TBPTT
@@
$\omega:\mathcal{U}(15,85)$, $b^{\prime}:\mathcal{U}(0.1,1),$
@@
$\omega:\mathcal{U}(10,50)$, $b^{\prime}:\mathcal{U}(1,6)$,
@@
ECG-QT BRF-RSNN BHRF-RSNN ALIF-RSNN
@@
$\omega:\mathcal{U}(3,5)$, $b^{\prime}:\mathcal{U}(0.1,1)$,
@@
$\omega:\mathcal{U}(7,11)$, $b^{\prime}:\mathcal{U}(0.1,1)$,
@@
SHD BRF-RSNN BHRF-RSNN ALIF-RSNN
@@
$\omega:\mathcal{U}(5,10)$, $b^{\prime}:\mathcal{U}(2,3)$,
@@
$\omega:\mathcal{U}(3,10)$, $b^{\prime}:\mathcal{U}(0.1,1)$,
@@
A.13 Learning speed-up with forward gradient injection (FGI) and TorchScript
As a further exploration, we apply forward gradient injection (FGI) from Otte (2024) for modeling the surrogate gradient functions. Let $x$ be a computation node, $f$ a non-differentiable function (e.g. step function), $g^{\prime}$ a function that we want to use as derivative for $f$, and $\operatorname{sg}$ the stop gradient operator. Instead of overriding the backward() method, we can just write:
$h$ $=x\cdot\operatorname{sg}({g^{\prime}(x)})$ (44)
$y$ $=h-\operatorname{sg}(h)+\operatorname{sg}(f(x))$ (45)
With FGI + TorchScript we achieve speed-ups by a factor of more than two for all dataset as shown in Table 4, while maintaining performance.
Table 4: BRF-RSNN training speed up with model optimization methods over standard backward() baseline.
Optimization Gradient S-MNIST PS-MNIST ECG SHD
Time (hrs) Ratio Time (hrs) Ratio Time (hrs) Ratio Time (min) Ratio
- backward() 23.8 1×1$\times$1 × 31.0 1×1$\times$1 × 8.5 1×1$\times$1 × 40.9 1×1$\times$1 ×
TorchScript backward() 11.3 2.1×2.1$\times$2.1 × 13.8 $\times2$ 4.1 2.1×2.1$\times$2.1 × 16.4 2.5×2.5$\times$2.5 ×
TorchScript FGI 10.2 $\mathbf{2.3}\times$ 12.8 $\mathbf{2.4}\times$ 3.5 $\mathbf{2.4}\times$ 14.0 $\mathbf{2.9}\times$